# Sampling by averaging: A multiscale approach to score estimation

### v1

@@ C.3 Bias of the annealed Langevin dynamics
$\begin{cases}\mathrm{d}\bar{Z}_{t}=\left(M^{-1}\bar{U}_{t}+\hat{v}_{1,t}(\bar{Z}_{t},\bar{U}_{t})\right)\mathrm{d}t\\ \mathrm{d}\bar{U}_{t}=\left(\nabla\log\hat{\mu}_{t}(\bar{Z}_{t})-\Gamma M^{-1}\bar{U}_{t}+\hat{v}_{2,t}(\bar{Z}_{t},\bar{U}_{t})\right)\mathrm{d}t+\sqrt{2\Gamma}\mathrm{d}B_{t},\end{cases}\quad\quad\quad\quad\text{for $t\in[0,1/\kappa]$. }$
@@
$=-\nabla_{x}\cdot\left(\hat{p}_{t}\hat{v}_{1,t}\right)-\nabla_{v}\cdot\left(\hat{p}_{t}\hat{v}_{2,t}\right),\quad t\in[0,1/\kappa].$
@@ Appendix D Experiment details
The code to reproduce our experiments is available at https://anonymous.4open.science/r/sampling_by_averaging-41F2. All experiments were implemented using JAX.
@@ D.1 Target distributions
The mixture of Student’s t distributions consists of 10 standard Student’s t distributions, each with 2 degrees of freedom ($t_{2}$). Following [Chen et al., 2025; Blessing et al., 2024], the mean of each component is sampled uniformly from a hypercube with a side length of 10. We evaluate this benchmark in dimensions 2, 10 and 50.
@@
where $\sigma$ denotes the sigmoid function. Following [Grenioux et al., 2024], we place independent Gaussian priors on the parameters $p(w,b)=\mathcal{N}(w;0,I_{d})\mathcal{N}(x;0,2.5^{2})$.
@@
The goal is to sample metastable states of the stochastic Allen-Cahn equation $\phi^{4}$ model (dimension 100) [Albergo et al., 2019; Gabrié et al., 2022]. The $\phi^{4}$ model is a continuous relaxation of the Ising model which is used to study phase transitions in statistical mechanics. Following [Albergo et al., 2019; Gabrié et al., 2022; Grenioux et al., 2024], we consider a version of the model discretised on a 1-dimensional grid of size $d=100$. Each configuration of the model is represented by a $d$-dimensional vector $(\phi_{i})_{i=1}^{d}$. We clip the field to 0 at the boundaries by defining $\phi_{0}=\phi_{d+1}=0$. The negative log-density of the distribution is defined as
@@ D.3 Algorithms and hyperparameters
For each baseline algorithm, we perform a grid search over a predefined set of hyperparameter values. Selection is based on the corresponding performance metric, computed using 4096 samples across 30 random seeds to account for variability and ensure robustness. The selected hyperparameters for each algorithm, together with the tuning grids, are summarised below.
@@
    $T$ $0.30$ $0.30$ $0.35$ $0.10$ $0.10$
  • For the SLIPS algorithm, we adopt the implementation and hyperparameters provided in the original paper [Grenioux et al., 2024]. To ensure a fair comparison with other algorithms initialised from a standard Gaussian distribution, we set the scalar variance parameter $R_{\pi}=\sqrt{d}$, yielding $\sigma=R_{\pi}/\sqrt{d}=1$. While this choice may degrade SLIPS’ performance compared to using an optimally tuned $R_{\pi}$, estimating such a parameter in practical scenarios is often non trivial. Nevertheless, for completeness, we report below the performance of SLIPS with its optimal parameters alongside our algorithms. In this setting, the results are comparable. However, a key advantage of our method is that it does not require estimation of the scalar variance. Moreover, if we initialise our algorithms using the same informed choice $R_{\pi}$ by setting the base distribution $\nu\sim\mathcal{N}(0,\sigma^{2}I)$, with $\sigma=R_{\pi}/\sqrt{d}$, we observe a performance improvement—particularly in the overdamped regime.
@@ D.4 Computation time
    All experiments were conducted on a GPU server consisting of eight Nvidia GTX 3090 Ti GPU cards, 896 GB of memory and 14TB of local on-server data storage. Each GPU has 10496 cores as well as 24 GB of memory.
@@ E.1 Comparison of overdamped and underdamped dynamics
As noted in the main text, the overdamped version of MultALMC requires a small value of $\kappa$, which corresponds to a slowly varying dynamics driven by $\nabla\log\hat{\mu}_{t}$ to perform well in practice. However, this leads to a large number of discretisation steps since we use a small step size with respect to this slowly changing dynamics resulting in high computational costs. In contrast, the underdamped version achieves better performance without the need for such small step sizes, thanks to the faster convergence properties of underdamped dynamics [Eberle and Lörler, 2024]. To further analyse this, we compare in Table 10 the performance and number of energy evaluations of the overdamped and underdamped versions of MultALMC, as specified in Algorithms 1 and 2, respectively, on the 40-MoG benchmark across different dimensions. The results show that the overdamped version requires approximately an order of magnitude more energy evaluations (and hence time steps) to achieve performance comparable to the underdamped version.
Table 10: Performance metrics on the 40-MoG benchmark across varying dimensions, averaged over 30 runs with different random seeds, along with the number of energy evaluations for the overdamped and underdamped versions of MultALMC, as defined in Algorithms 1 and 2, respectively.
@@ E.2 Analysis of the impact of the number of SROCK steps in the discretisation
The number of SROCK steps used in our algorithms—MultALMC and MultCDiff—as implemented in Algorithms 1, 2, and 3, plays a critical role in balancing computational efficiency and numerical performance. Increasing the number of inner SROCK steps generally leads to better performance [Abdulle et al., 2018]; however, this comes at the cost of greater computational overhead. Ideally, we aim to identify the minimum number of steps required to maintain strong performance and computational efficiency.
@@
Table 11: Performance on the 40-MoG benchmark in 50 dimensions using varying numbers of SROCK steps in the discretisation of MultALMC and MultCDiff, as proposed in Algorithms 2 and 3, respectively. The number of sampling steps $L$ for the slow process is held fixed. The standard deviation is computed over 30 runs with different random seeds. The regularisation parameter for the entropy regularised Wasserstein-2 distance $W_{2}^{\delta}$ is set to $\delta=0.05$.
@@ E.3 Stability across hyperparameter values
Our algorithms demonstrate robustness and stability across a broad range of hyperparameter values, as shown in tables 12-15. These results suggest that while some tuning is beneficial, our methods do not require highly sensitive or exhaustive hyperparameter optimisation.
Table 12: Ablation results on the 40-component Mixture of Gaussians benchmark in 50 dimensions for the scale separation parameter $\varepsilon$. Results are averaged over 30 random seeds.
$\mathbf{0.01}$ $\mathbf{0.03}$ $\mathbf{0.05}$ $\mathbf{0.07}$ $\mathbf{0.10}$ $\mathbf{0.12}$ $\mathbf{0.15}$
[…]
Table 13: Ablation results on the 40-component Mixture of Gaussians benchmark in 50 dimensions for the number of steps. Results are averaged over 30 random seeds.
Steps $\mathbf{5\times 10^{5}}$ $\mathbf{7.5\times 10^{5}}$ $\mathbf{1\times 10^{6}}$ $\mathbf{2.5\times 10^{6}}$ $\mathbf{5\times 10^{6}}$
[…]
Table 14: Ablation results on the Sonar benchmark for the scale separation parameter $\varepsilon$. Results are averaged over 30 random seeds. The performance metric is the average predictive posterior log-likelihood on a test dataset.
$\varepsilon$ $\mathbf{0.01}$ $\mathbf{0.03}$ $\mathbf{0.05}$ $\mathbf{0.07}$ $\mathbf{0.10}$ $\mathbf{0.12}$ $\mathbf{0.15}$
[…]
Table 15: Ablation results on the Sonar benchmark for the number of steps. Results are averaged over 30 random seeds. The performance metric is the average predictive posterior log-likelihood on a test dataset.
Steps $\mathbf{5\times 10^{5}}$ $\mathbf{7.5\times 10^{5}}$ $\mathbf{1\times 10^{6}}$ $\mathbf{2.5\times 10^{6}}$ $\mathbf{5\times 10^{6}}$
[…]
@@ Appendix F Limitations and future work
First, the theoretical guarantees presented in Section 4 rely on stringent assumptions, such as strong convexity of the potential $V_{\pi}$ and Lipschitz continuity of the score function of the target $\nabla\log\pi$. Relaxing these assumptions—e.g., to strong convexity outside a compact region or to satisfying weak functional inequalities—could broaden the applicability of our analysis. The former would accommodate target distributions that are multimodal within a compact region and Gaussian-like in the tails, while the latter could capture heavy-tailed distributions.

### v2

@@ C.3 Bias of the annealed Langevin dynamics
$\begin{cases}\mathrm{d}\bar{Z}_{t}=\left(M^{-1}\bar{U}_{t}+\hat{v}_{1,t}(\bar{Z}_{t},\bar{U}_{t})\right)\mathrm{d}t\\ \mathrm{d}\bar{U}_{t}=\left(\nabla\log\hat{\mu}_{t}(\bar{Z}_{t})-\Gamma M^{-1}\bar{U}_{t}+\hat{v}_{2,t}(\bar{Z}_{t},\bar{U}_{t})\right)\mathrm{d}t+\sqrt{2\Gamma}\mathrm{d}B_{t},\end{cases}\qquad\qquad\qquad\qquad\text{for $t\in[0,1/\kappa]$. }$
@@
$=-\nabla_{x}\cdot\left(\hat{p}_{t}\hat{v}_{1,t}\right)-\nabla_{v}\cdot\left(\hat{p}_{t}\hat{v}_{2,t}\right),\qquad t\in[0,1/\kappa].$
@@ Appendix D Experiment details
The code to reproduce our experiments is available at https://github.com/paulaoak/sampling_by_averaging.git. All experiments were implemented using JAX.
@@ D.1 Target distributions
The mixture of Student’s t distributions consists of 10 standard Student’s t distributions, each with 2 degrees of freedom ($t_{2}$). Following [Chen et al., 2025, Blessing et al., 2024], the mean of each component is sampled uniformly from a hypercube with a side length of 10. We evaluate this benchmark in dimensions 2, 10 and 50.
@@
where $\sigma$ denotes the sigmoid function. Following [Grenioux et al., 2024], we place independent Gaussian priors on the parameters $p(w,b)=\mathcal{N}(w;0,I_{d})\mathcal{N}(b;0,2.5^{2})$.
@@
The goal is to sample metastable states of the stochastic Allen-Cahn equation $\phi^{4}$ model (dimension 100) [Albergo et al., 2019, Gabrié et al., 2022]. The $\phi^{4}$ model is a continuous relaxation of the Ising model which is used to study phase transitions in statistical mechanics. Following [Albergo et al., 2019, Gabrié et al., 2022, Grenioux et al., 2024], we consider a version of the model discretised on a 1-dimensional grid of size $d=100$. Each configuration of the model is represented by a $d$-dimensional vector $(\phi_{i})_{i=1}^{d}$. We clip the field to 0 at the boundaries by defining $\phi_{0}=\phi_{d+1}=0$. The negative log-density of the distribution is defined as
@@ D.3 Algorithms and hyperparameters
For each baseline algorithm, we perform a grid search over a predefined set of hyperparameter values. Selection is based on the corresponding performance metric, computed using 4096 samples. The selected hyperparameters for each algorithm are summarised below.
@@
    $T$ $\log(0.70)$ $\log(0.70)$ $\log(0.75)$ $\log(0.95)$ $\log(0.95)$
  • For the SLIPS algorithm, we adopt the implementation and hyperparameters provided in the original paper [Grenioux et al., 2024]. To ensure a fair comparison with other algorithms initialised from a standard Gaussian distribution, we set the scalar variance parameter $R_{\pi}=\sqrt{d}$, yielding $\sigma=R_{\pi}/\sqrt{d}=1$. While this choice may degrade SLIPS’ performance compared to using an optimally tuned $R_{\pi}$, estimating such a parameter in practical scenarios is often non trivial. Nevertheless, for completeness, we report below the performance of SLIPS with its optimal parameters alongside our algorithms. In this setting, the results are comparable. However, a key advantage of our method is that it does not require estimation of the scalar variance. Moreover, if we initialise our algorithms using the same informed choice $R_{\pi}$ by setting the base distribution $\nu\sim\mathcal{N}(0,\sigma^{2}I)$, with $\sigma=R_{\pi}/\sqrt{d}$, we observe a performance improvement, particularly in the overdamped regime.
@@ D.4 Computation time
    All experiments were conducted on a GPU server consisting of eight Nvidia GeForce RTX 3090 Ti GPU cards, 896 GB of memory and 14TB of local on-server data storage. Each GPU has 10496 cores as well as 24 GB of memory.
@@ E.1 Comparison of overdamped and underdamped dynamics
As noted in the main text, the overdamped version of MultALMC requires a small value of $\kappa$, which corresponds to a slowly varying dynamics driven by $\nabla\log\hat{\mu}_{t}$ to perform well in practice. However, this leads to a large number of discretisation steps since we use a small step size with respect to this slowly changing dynamics resulting in high computational costs. In contrast, the underdamped version achieves better performance without the need for such small step sizes, thanks to the faster convergence properties of underdamped dynamics [Eberle and Lörler, 2024]. To further analyse this, we compare in Table 10 the performance and number of energy evaluations of the overdamped and underdamped versions of MultALMC, as specified in Algorithms 2 and 3, respectively, on the 40-MoG benchmark across different dimensions. The results show that the overdamped version requires approximately an order of magnitude more energy evaluations (and hence time steps) to achieve performance comparable to the underdamped version.
Table 10: Performance metrics on the 40-MoG benchmark across varying dimensions, averaged over 30 runs with different random seeds, along with the number of energy evaluations for the overdamped and underdamped versions of MultALMC, as defined in Algorithms 2 and 3, respectively.
@@ E.2 Analysis of the impact of the number of SROCK steps in the discretisation
The number of SROCK steps used in our algorithms—MultALMC and MultCDiff—as implemented in Algorithms 2, 3, and 4, plays a critical role in balancing computational efficiency and numerical performance. Increasing the number of inner SROCK steps generally leads to better performance [Abdulle et al., 2018]; however, this comes at the cost of greater computational overhead. Ideally, we aim to identify the minimum number of steps required to maintain strong performance and computational efficiency.
@@
Table 11: Performance on the 40-MoG benchmark in 50 dimensions using varying numbers of SROCK steps in the discretisation of MultALMC and MultCDiff, as proposed in Algorithms 3 and 4, respectively. The number of sampling steps $L$ for the slow process is held fixed. The standard deviation is computed over 30 runs with different random seeds. The regularisation parameter for the entropy regularised Wasserstein-2 distance $W_{2}^{\delta}$ is set to $\delta=0.05$.
@@ E.3 Stability across hyperparameter values
Our algorithms demonstrate robustness and stability across a broad range of hyperparameter values, as shown in Figures 3 and 4. These results suggest that while some tuning is beneficial, our methods do not require highly sensitive or exhaustive hyperparameter optimisation.
(a) Performance against scale separation parameter $\varepsilon$.
(b) Performance against number of steps.
Figure 3: Ablation results on the 40-component Mixture of Gaussians benchmark in 50 dimensions. Results are averaged over 30 random seeds.
(a) Performance against scale separation parameter $\varepsilon$.
(b) Performance against number of steps.
Figure 4: Ablation results on the Sonar benchmark. Results are averaged over 30 random seeds. The performance metric is the average predictive posterior log-likelihood on a test dataset.
@@ Appendix F Limitations and future work
First, the theoretical guarantees presented in Section 4 rely on stringent assumptions, such as strong convexity of the potential $V_{\pi}$ and Lipschitz continuity of the score function of the target $\nabla\log\pi$. Relaxing these assumptions, e.g., to strong convexity outside a compact region or to satisfying weak functional inequalities, could broaden the applicability of our analysis. The former would accommodate target distributions that are multimodal within a compact region and Gaussian-like in the tails, while the latter could capture heavy-tailed distributions.